OpenStudy (kdhdjhio):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 2i

jimthompson5910 (jim_thompson5910):

how far did you get with this one @kdhdjhio ?

OpenStudy (kdhdjhio):

I can't figure this out along with like 4 others, I've tried but I just don't understand.

jimthompson5910 (jim_thompson5910):

Do you agree that if x = 5 is a root of f(x), then x-5 is a factor of f(x)?

OpenStudy (kdhdjhio):

I think so

jimthompson5910 (jim_thompson5910):

if x = -3 is a root, then the factor is ________ (fill in the blank)

jimthompson5910 (jim_thompson5910):

hint: start with x = -3 and try to get everything to the left side. The right side should have 0

OpenStudy (kdhdjhio):

Ugh..I don't understand

jimthompson5910 (jim_thompson5910):

x = -3 x+3 = -3+3 ... add 3 to both sides x+3 = 0 so if x = -3 is a root, then x+3 is a factor

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (kdhdjhio):

Yes, that makes more sense

jimthompson5910 (jim_thompson5910):

ok so if x = 5 is a root, then x-5 is a factor if x = -3 is a root, then x+3 is a factor so far we have the factors (x-5) and (x+3) what do you get when you multiply out (x-5)*(x+3) ?

jimthompson5910 (jim_thompson5910):

you can use FOIL or the box method. I prefer the box method

jimthompson5910 (jim_thompson5910):

box method